Haunted Forest, Inc.is selling fog machines. Use the following information about Haunted Forest, Inc. to answer the following questions. Average selling price per unit $337. Variable cost per unit $214 Units sold 399 Fixed costs $17,392 Interest expense $4,938 Based on the data above, what will be the resulting percentage change in earnings per share if they expect units produced and sold to change -4.4 percent?

Round the answer to two decimals
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

B C D E F Unit 399 399 Sales Variable cost Fixed Cost EBIT Interest Expeses EBT rate Amount $337.00 $134,463.00 $214.00 $85,3

Cell reference -

A B C D E F Unit 399 399 G H I After change in Sale unit Unit rate Amount =C3*(1+C12) 337 =G3*H3 =G3 214 =G4*H4 rate Amount 3

Please note:

Degree of financial leverage = EBIT/EBT

also,

Degree of finanacial leverage = % change in EPS/% change in EBIT

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.